# Access Hoare Logic

### v2

@@ 7. Conclusion
In this brief note, we made the point that access security cannot be addressed naturally by standard Hoare logic. To overcome this issue, we introduced access Hoare logic where the reasoning is in the reverse direction compared to Hoare logic, namely from postconditions to preconditions. We gave three examples to demonstrate our point, one regarding electronic keys, a second considering the cryptocurrency Bitcoin, and a third about granting access involving a key and a list of keys. We introduced rules for access Hoare logic, and showed that these rules are sound and complete. We provided a link between access Hoare logic and Hoare logic by showing that weakest preconditions for Hoare logic coincides with strongest preconditions for access Hoare logic for total programs.
While access Hoare logic can be reduced to Hoare logic by introducing negation, we argued that this would have several negative consequences, including breaking verifiability using intuitionistic theorem provers — indeed, careful inspection shows that all proofs in this paper (except for Theorem 3.1, which needs to be replaced by Remark 3.2) only use intuitionistic logic, which we have confirmed using Agda (agda:Documentation) which is based on intuitionistic type theory. Details of that formalization will appear in a separate publication.
There are, of course, many areas that we did not touch upon. Basically, everything that has been investigated for Hoare logic should now be revisited in the context of access Hoare logic. Instead of aiming for a complete list, we mention two obvious areas. In our approach to soundness and completeness, we define the semantics of expressions and consequences in an informal way. There is a vast literature on ways to make this more precise, by considering formal semantics, their axiomatization and implications on Hoare logic (apt:1981). As a second area, the relationship to Dijkstra’s predicate transformer semantics and its variations need to be clarified (dijkstra:1975). Finally, detailed case studies are needed to demonstrate the usefulness of access Hoare logic for access security.
In connection to this paper, the authors have filed a patent entitled Verifying Access Security of a Computer Program (beckmannSetzer:Patent:AccessHoareLogic:FilingUK). This patent utilizes verification condition generation for access Hoare logic, which will be detailed in a separate publication.
@@ Acknowledgements.
We would like to thank Fahad F. Alhabardi. This article builds upon and expands the use of weakest preconditions for access security, which were explored during his PhD research (alhabardi:PhD:SmartContractsAgda). We would also like to thank John V Tucker for valuable comments on an earlier draft of this paper.

### v3

@@ 9. Conclusion
In this note, we made the point that access security cannot be addressed naturally by standard Hoare logic. To overcome this issue, we introduced access Hoare logic where the reasoning is in the reverse direction compared to Hoare logic, namely from post-conditions to pre-conditions. We gave three examples to demonstrate our point, one regarding electronic keys, a second considering the cryptocurrency Bitcoin, and a third about granting access involving a key and a list of keys. We introduced rules for access Hoare logic, and showed that these rules are sound and complete. We provided a link between access Hoare logic and Hoare logic by showing that weakest pre-conditions for Hoare logic coincides with strongest pre-conditions for access Hoare logic for total programs. We also demonstrated a fundamental difference between access Hoare logic and incorrectness logic and related approaches, in that access Hoare logic provides a complementary view to standard Hoare logic in the context of backwards transformers, while incorrectness logic does this for forward transformers.
While access Hoare logic can be reduced to Hoare logic by introducing negation, we argued that this would have several negative consequences, including breaking verifiability using intuitionistic theorem provers — indeed, careful inspection shows that all proofs in this paper (except for Theorem 3.1, which needs to be replaced by Remark 3.2) only use intuitionistic logic, which we have confirmed using Agda (Agda Community, 2025), an interactive theorem prover based on intuitionistic type theory. Details of that formalization will appear in a separate publication.
There are, of course, many areas that we did not touch upon. Basically, everything that has been investigated for Hoare logic should now be revisited in the context of access Hoare logic. Instead of aiming for a complete list, we mention two obvious areas. In our approach to soundness and completeness, we take assertions to be semantical. There is a vast literature on ways to make this more precise, by fixing the syntax of the expressions in the language, considering formal semantics, their axiomatization and implications on Hoare logic (Apt, 1981). As a second area, the relationship to Dijkstra’s predicate transformer semantics and its variations need to be clarified (Dijkstra, 1975). Finally, detailed case studies are needed to demonstrate the usefulness of access Hoare logic for access security.
In connection to this paper, the authors have filed a patent entitled Verifying Access Security of a Computer Program (Beckmann and Setzer, 2025). This patent utilizes verification condition generation for access Hoare logic, which will be detailed in a separate publication.
@@ Acknowledgements.
We would like to thank Fahad F. Alhabardi. This article builds upon and expands the use of weakest pre-conditions for access security, which were explored during his PhD research (alhabardi:PhD:SmartContractsAgda). We would also like to thank John V Tucker for valuable comments on an earlier draft of this paper.